OpenStudy (anonymous):

y=√1+√1+√x find the derivative Absolutely no idea how to do this one.

OpenStudy (anonymous):

sqrt(1) and sqrt(1) are just constants, the derivatives of constants = 0

OpenStudy (bahrom7893):

1/(2sqrt(x))

OpenStudy (anonymous):

and dy/dx sqrt(x) =x^1/2 = 1/2(x)^-1/2 by the power rule

OpenStudy (anonymous):

The answer in the back of the book is 1 ------------------------------------------ 8√x (√1+√x)(√1+√1+√x)

myininaya (myininaya):

\[y=\sqrt{1+\sqrt{1+\sqrt{x}}}\] is this what you meant ?

OpenStudy (anonymous):

ah yes that's it lol sorry if that makes it completely different.

myininaya (myininaya):

yes this makes it completely different

myininaya (myininaya):

You need to know chain rule fo' sho'

OpenStudy (anonymous):

Ya I'm just not sure which is the inside and which is out side since there's 3 terms.

myininaya (myininaya):

\[y=(1+(1+x^\frac{1}{2})^\frac{1}{2})^\frac{1}{2}\] Ok first of all it might help you to look this first: \[y=f(g(h(x))) => y'=f'(g(h(x)) \cdot g'(h(x)) \cdot h'(x)\]

myininaya (myininaya):

\[y'=\frac{1}{2}(1+(1+x^\frac{1}{2})^\frac{1}{2})^{\frac{1}{2}-1} (1+(1+x^\frac{1}{2})^\frac{1}{2})'\] But what is \[(1+(1+x^\frac{1}{2})^\frac{1}{2})'\] \[=(0+\frac{1}{2}(1+x^\frac{1}{2})^{\frac{1}{2}-1})(1+x^\frac{1}{2})'\] But what is \[(1+x^\frac{1}{2})'\] \[=0+\frac{1}{2}x^{\frac{1}{2}-1}\] So we have.... \[y'=\frac{1}{2}(1+(1+x^\frac{1}{2})^\frac{1}{2})^\frac{-1}{2}(0+\frac{1}{2}(1+x^\frac{1}{2})^\frac{-1}{2})(0+\frac{1}{2}x^{\frac{-1}{2}})\]

myininaya (myininaya):

lol that is nasty honestly

OpenStudy (anonymous):

Ha ya I just hope that's not on my quiz tomorrow. Thanks for the help.

myininaya (myininaya):

hey i got something better looking that you might understand better you still there?

OpenStudy (anonymous):

That made a lot more sense. I appreciate the time you took to do that.

myininaya (myininaya):

I'm glad it looked better to you I hope you understand what I did

OpenStudy (anonymous):

I do indeed.